### v1

# Primordial Black Hole Formation in Rastall Gravity: Shifted Collapse Threshold and Exponential Abundance Sensitivity

Abstract
Primordial black holes formed in the early universe are compelling candidates for dark matter. We investigate their production in Rastall gravity, a modification of general relativity that introduces a non-minimal coupling between matter and geometry through the non-conservation of the energy-momentum tensor. Analyzing cosmological perturbations during radiation domination, we demonstrate that the Rastall parameter fundamentally alters the collapse dynamics, modifying the growth of density fluctuations, the critical threshold for black hole formation, and the fluctuation amplitude at horizon crossing. Current cosmological constraints from Big Bang Nucleosynthesis, the cosmic microwave background, and large-scale structure restrict the Rastall parameter to small values, yet within this allowed range PBH production can be altered by orders of magnitude compared to general relativity. Our results establish primordial black holes as novel probes of modified gravity.
Keywords: Primordial black holes, Rastall gravity, dark matter, cosmological perturbations
1 Introduction
Primordial black holes (PBHs) have emerged as compelling dark matter candidates and unique probes of the early universe [1, 2, 3, 4, 5, 6]. Formed from the gravitational collapse of large-amplitude density perturbations during radiation domination, PBHs bridge the gap between early universe physics and observational astronomy, particularly through gravitational wave astronomy [7, 8, 9, 10, 11, 12, 13]. The standard formation mechanism in general relativity (GR) requires specific enhancements of the primordial power spectrum, often necessitating fine-tuned inflationary dynamics [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
Modified gravity theories offer alternative pathways to achieve the conditions necessary for PBH formation [24, 25, 26, 27, 28]. Among these, Rastall gravity [29, 30] presents a unique and intriguing modification. Rastall’s proposal relaxes the covariant conservation of the energy-momentum tensor, postulating instead a non-minimal coupling between matter and geometry:
$\nabla_{\mu}T^{\mu\nu}=\lambda\nabla^{\nu}R,$ (1)
where $\lambda$ is the Rastall parameter and $R$ is the Ricci scalar. while preserving the Bianchi identity $\nabla_{\mu}G^{\mu\nu}=0$. This leads to the modified field equations
$G_{\mu\nu}=\kappa\left(T_{\mu\nu}-\frac{\lambda}{1-4\lambda}g_{\mu\nu}T\right),$ (2)
which reduce to general relativity (GR) in the limit $\lambda\to 0$. [31, 32, 33, 34, 35, 36, 37, 38].
A longstanding question in the literature concerns whether Rastall gravity represents a genuine modification of gravitational dynamics or merely a reformulation of GR with a redefined effective energy–momentum tensor. In particular, it has been emphasized that the field equations of Rastall gravity can formally be rewritten as Einstein equations with an effective stress tensor,
$G_{\mu\nu}=\kappa\,T^{\rm(eff)}_{\mu\nu}$, (3)
where
$T^{\rm(eff)}_{\mu\nu}=T_{\mu\nu}-\frac{\lambda}{1-4\lambda}g_{\mu\nu}T.$ (4)
In this sense, Rastall gravity may be interpreted as standard GR coupled to a modified matter sector with non-standard conservation properties[35].
The physical interpretation depends on how the matter sector is defined. If $T_{\mu\nu}$ represents a physically conserved matter tensor derived from a matter Lagrangian minimally coupled to the metric, then Rastall gravity effectively modifies the coupling between geometry and matter through the trace term. The modified conservation equation then implies that matter and geometry exchange energy–momentum in a manner not present in GR. From this perspective, Rastall gravity alters the dynamical response of matter perturbations to curvature and therefore leads to physically distinct cosmological perturbation dynamics.
In homogeneous cosmology with a radiation fluid ($T=0$), the background expansion coincides exactly with GR. However, perturbations probe variations of the curvature scalar and the trace sector, and the modified conservation equation introduces additional structure at the perturbative level. Therefore, even in scenarios where the background cosmology is indistinguishable from GR, linear perturbations can encode nontrivial dependence on the Rastall parameter $\lambda$.
In the present work, we adopt the interpretation that $T_{\mu\nu}$ represents the physical matter tensor of a radiation fluid satisfying the modified conservation law above. We do not treat the theory as a mere redefinition of matter variables within GR; instead, we take the modified conservation equation as defining the physical coupling between matter and geometry. Under this interpretation, deviations in the evolution of density perturbations and, consequently, in primordial black hole formation are physically meaningful predictions of the theory.
It is worth emphasizing that the observational distinguishability of Rastall gravity arises entirely from dynamical effects in perturbations rather than from additional propagating gravitational degrees of freedom. The theory does not introduce extra tensor polarizations at the linearized level around Minkowski space. Instead, its distinctive features emerge through modified matter–geometry coupling, which can lead to measurable consequences in early-universe structure formation processes such as primordial black hole production.
The Rastall framework has been extensively studied in various contexts. Black hole solutions in Rastall gravity have been investigated by several authors [39, 40, 41, 42, 43, 44, 45, 46], revealing interesting modifications to horizon structures and thermodynamic properties. Cosmological implications have been explored in works examining background evolution [47, 48, 49, 50, 51, 52, 53, 54], perturbation dynamics [55, 56, 57, 58, 59, 60, 61, 62], and structure formation [63, 64, 65]. Gravitational wave propagation in Rastall gravity has also been analyzed [55, 56, 59, 60, 61], showing modifications to the evolution of perturbation parameters.
2 Rastall Gravity: Fundamentals and Background Evolution
The Rastall field equations are derived by maintaining the Bianchi identity $\nabla^{\mu}G_{\mu\nu}=0$ while incorporating the non-conservation postulate (1). This leads to:
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\kappa\left(T_{\mu\nu}-\frac{\lambda}{1-4\lambda}g_{\mu\nu}T\right),$ (5)
where $\kappa=8\pi G$ and $T=T^{\mu}_{\mu}$ is the trace of the energy-momentum tensor. The parameter $\lambda$ must satisfy $\lambda\neq 1/4$ to avoid singular behavior. For $\lambda=0$, GR is recovered. The coupling between geometry and matter is evident through the trace-dependent term.
For a perfect fluid with equation of state $p=w\rho$, the energy-momentum tensor and its trace are:
$T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu},\quad T=-\rho+3p=(3w-1)\rho.$ (6)
Considering a flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric:
$ds^{2}=-dt^{2}+a^{2}(t)\delta_{ij}dx^{i}dx^{j},$ (7)
where $a(t)$ is the scale factor and $H=\dot{a}/a$ is the Hubble parameter.
Substituting into the field equations (5) yields the modified Friedmann equations:
$3H^{2}$ $=\kappa\left[1-\frac{\lambda(1-3w)}{1-4\lambda}\right]\rho,$ (8)
$2\dot{H}+3H^{2}$ $=-\kappa\left[w-\frac{\lambda(1-3w)}{1-4\lambda}\right]\rho.$ (9)
The conservation law (1) for the FLRW metric gives the time component:
$\dot{\rho}+3H(\rho+p)=\lambda\dot{R},\quad R=6(2H^{2}+\dot{H}).$ (10)
For radiation domination, $w=1/3$, so 13w=01-3w=0. Equations (8)-(9) simplify dramatically:
$3H^{2}$ $=\kappa\rho,$ (11)
$2\dot{H}+3H^{2}$ $=-\frac{\kappa\rho}{3}.$ (12)
Thus, the background Friedmann equations are identical to those of general relativity. From these, we recover the standard relations:
$H^{2}$ $=\frac{8\pi G}{3}\rho,$ (13)
$\dot{H}$ $=-2H^{2},$ (14)
$\rho$ $\propto a^{-4},\quad a\propto t^{1/2}\propto\tau,$ (15)
where $\tau$ is conformal time defined by $d\tau=dt/a$.
Now examine the conservation equation (10) for radiation. Using $p=\rho/3$:
$\dot{\rho}+4H\rho=\lambda\dot{R}.$ (16)
From the Friedmann equations, we compute the Ricci scalar:
$R=6(2H^{2}+\dot{H})=6(2H^{2}-2H^{2})=0.$ (17)
Since $R=0$, we have $\dot{R}=0$, and (16) reduces to the standard conservation law $\dot{\rho}+4H\rho=0$. Consequently, the background evolution for a single radiation fluid is indistinguishable from GR.
While the background dynamics for pure radiation are identical to GR, differences emerge at the perturbation level. This makes PBH formation during radiation domination an ideal probe of Rastall gravity, as any deviations from GR must arise from perturbations rather than background expansion.
3 Scalar Perturbations and Density Contrast Evolution
We work in the Newtonian (longitudinal) gauge with $\Phi=\Psi$ (no anisotropic stress):
$ds^{2}=-(1+2\Phi)dt^{2}+a^{2}(t)(1-2\Phi)\delta_{ij}dx^{i}dx^{j}.$ (18)
The perturbed energy-momentum tensor components for a perfect fluid are:
$\delta T^{0}_{0}$ $=-\delta\rho,$ (19)
$\delta T^{0}_{i}$ $=(\rho+p)\partial_{i}v,$ (20)
$\delta T^{i}_{j}$ $=\delta p\,\delta^{i}_{j},$ (21)
where $\delta\rho$ is the density perturbation, $\delta p=c_{s}^{2}\delta\rho$ is the pressure perturbation (with $c_{s}^{2}=w$ for adiabatic perturbations), and $v$ is the velocity potential.
The perturbed Rastall field equations are obtained by varying (5). The 00-component in Fourier space ($\nabla^{2}\rightarrow-k^{2}$) yields:
$\frac{2k^{2}}{a^{2}}\Phi+6H(\dot{\Phi}+H\Phi)=\kappa\left[\left(1-\frac{4\lambda}{1-4\lambda}c_{s}^{2}\right)\delta\rho\right].$ (22)
The 0i-component gives:
$2\partial_{i}(\dot{\Phi}+H\Phi)=\kappa(\rho+p)\partial_{i}v.$ (23)
The perturbed Rastall conservation law $\delta(\nabla_{\mu}T^{\mu\nu})=\lambda\nabla^{\nu}\delta R$ provides two crucial equations. The time component ($\nu=0$) is:
$\delta(\nabla_{\mu}T^{\mu 0})=-\lambda\partial_{0}\delta R.$ (24)
The left-hand side is the perturbed energy conservation equation:
$\delta(\nabla_{\mu}T^{\mu 0})=\dot{\delta\rho}+3H(\delta\rho+\delta p)+(\rho+p)\left(3\dot{\Phi}-\frac{k^{2}}{a^{2}}v\right).$ (25)
The spatial component ($\nu=i$) yields:
$\delta(\nabla_{\mu}T^{\mu i})=\frac{\lambda}{a^{2}}\partial^{i}\delta R.$ (26)
The left-hand side gives the perturbed momentum conservation:
$\delta(\nabla_{\mu}T^{\mu i})=(\rho+p)(\dot{v}+4Hv)+\partial^{i}\delta p+(\rho+p)\partial^{i}\Phi.$ (27)
The perturbed Ricci scalar to first order in Fourier space is:
$\delta R=\frac{2k^{2}}{a^{2}}\Phi+6\ddot{\Phi}+12H\dot{\Phi}+6(2H^{2}+\dot{H})\Phi.$ (28)
4 Derivation of the Master Equation for Density Contrast
We now combine these equations to derive a second-order equation for the density contrast $\delta\equiv\delta\rho/\rho$.
From the 0i-component (23):
$v=\frac{2(\dot{\Phi}+H\Phi)}{\kappa(\rho+p)}.$ (29)
Now, from the 00-component (22), for radiation with $c_{s}^{2}=1/3$ and working to linear order in $\lambda$ (the observationally allowed regime, $|\lambda|\ll 1$), we obtain:
$\delta\rho\approx-\frac{1}{\kappa}\left(1+\frac{4\lambda}{3}\right)\left(\frac{2k^{2}}{a^{2}}\Phi+6H(\dot{\Phi}+H\Phi)\right).$ (30)
Using $\delta=\delta\rho/\rho$ and the background relation $\kappa\rho=3H^{2}$ from (11), we get:
$\delta=-\frac{3H^{2}}{\kappa\rho}\left(1+\frac{4\lambda}{3}\right)\left(\frac{2k^{2}}{a^{2}}\Phi+6H(\dot{\Phi}+H\Phi)\right)^{-1}\ldots$ (31)
This is inverted to express $\Phi$ in terms of $\delta$. To linear order in perturbations and $\lambda$:
$\Phi\approx-\frac{a^{2}}{2k^{2}}\left[\frac{\kappa\rho}{3H^{2}}\left(1-\frac{4\lambda}{3}\right)\delta+6H(\dot{\Phi}+H\Phi)\right].$ (32)
This relation allows us to eliminate $\Phi$ in favor of $\delta$.
Inserting (29), (30), and the perturbed conservation laws. The momentum conservation (27) with (26) gives:
$(\rho+p)(\dot{v}+4Hv)+\partial^{i}\delta p+(\rho+p)\partial^{i}\Phi=\frac{\lambda}{a^{2}}\partial^{i}\delta R.$ (33)
Using the background relations for radiation domination:
  • $\rho+p=4\rho/3$
  • $\dot{H}=-2H^{2}$
  • $c_{s}^{2}=1/3$
  • $a\propto\tau^{1/2}$ in cosmic time, but we’ll switch to conformal time next.
Define conformal time $\tau$ with $d\tau=dt/a$, so that during radiation domination:
  • $a\propto\tau$
  • $\mathcal{H}=a^{\prime}/a=1/\tau$ (where $^{\prime}=d/d\tau$)
  • The comoving wavenumber $k$ is constant
Then keeping terms to linear order in $\lambda$, we obtain an equation for $\Phi$:
$\Phi^{\prime\prime}+\frac{4}{\tau}\Phi^{\prime}+\frac{k^{2}}{3}\Phi=\lambda\left(\frac{6}{\tau^{2}}\Phi^{\prime\prime}+\frac{24}{\tau^{3}}\Phi^{\prime}-\frac{4k^{2}}{3\tau^{2}}\Phi\right)+\mathcal{O}(\lambda^{2}).$ (34)
Using the relation between $\delta$ and $\Phi$ derived from (30) in conformal time:
$\delta\approx-\frac{2}{3}\left(1+\frac{4\lambda}{3}\right)\left(\frac{k^{2}\tau^{2}}{a^{2}}\Phi+3\tau\Phi^{\prime}+3\Phi\right).$ (35)
We differentiate (35) twice and substitute into (34). After some simplification, we arrive at the master equation for the density contrast in cosmic time:
$\boxed{\ddot{\delta}+2H\dot{\delta}-\left[\frac{3}{2}H^{2}-\frac{k^{2}}{3a^{2}}+\lambda\left(6H^{2}-\frac{4k^{2}}{3a^{2}}\right)\right]\delta=0.}$ (36)
This is our first key result. Compared to the GR equation, equation (36) reveals modifications to both the driving term and the effective sound speed.
4.1 Solution During Radiation Domination
[…]
$\frac{d^{2}\delta}{dx^{2}}+\frac{2}{x}\frac{d\delta}{dx}+\left[\frac{1}{3}-\frac{3}{2x^{2}}-\frac{4\lambda}{3}\left(1-\frac{3}{x^{2}}\right)\right]\delta=0.$ (37)
4.1.1 Super-horizon solution ($x\ll 1$)
For modes well outside the horizon, we seek a series solution of the form $\delta(x)=\delta_{0}(1+\alpha x^{2}+\mathcal{O}(x^{4}))$. Substituting and matching coefficients yields:
$\delta(x)\approx\delta_{0}\left[1+\left(\frac{3}{8}-\frac{2\lambda}{3}\right)x^{2}+\mathcal{O}(x^{4})\right].$ (38)
The $\lambda$-dependent term represents a modification to the growth rate of super-horizon modes.
4.1.2 Horizon crossing ($x=1$)
At horizon crossing, matching the super-horizon solution to the numerical solution gives the amplitude:
$\delta_{H}(k)=\frac{4}{9}(1+\alpha_{H}\lambda)\mathcal{P}_{\mathcal{R}}^{1/2}(k),$ (39)
where $\alpha_{H}\approx 0.3$ is a numerical coefficient derived from matching, and $\mathcal{P}_{\mathcal{R}}(k)$ is the primordial curvature power spectrum.
5 PBH Formation Criterion in Rastall Gravity
5.1 Critical Density Threshold for Collapse
The critical density contrast $\delta_{c}$ for PBH formation determines when an overdense region collapses against pressure forces. In general relativity, extensive numerical simulations [66, 67] have established $\delta_{c}^{\text{GR}}\approx 0.414$ for radiation domination.
In Rastall gravity, the threshold changes due to altered dynamics. We derive $\delta_{c}(\lambda)$ using two complementary approaches, keeping only terms linear in $\lambda$ for consistency with observational constraints.
5.1.1 Approach 1: Jeans scale modification
Figure 1: Critical density threshold $\delta_{c}(\lambda)$ as a function of the Rastall parameter $\lambda$. The solid line shows GR case, while the shaded region indicates the range of variation for different perturbation profiles based on numerical simulations.
The Jeans length $\lambda_{J}$ determines the scale below which pressure prevents collapse. From the dispersion relation derived from (36), we write $\omega^{2}=c_{s}^{2}k^{2}/a^{2}-4\pi G_{\text{eff}}\rho$, where $G_{\text{eff}}$ is the effective gravitational constant modified by Rastall gravity.
From (36), comparing with the standard form $\ddot{\delta}+2H\dot{\delta}-(4\pi G_{\text{eff}}\rho-c_{s}^{2}k^{2}/a^{2})\delta=0$, we identify:
$4\pi G_{\text{eff}}\rho$ $=\frac{3}{2}H^{2}-6\lambda H^{2}=\frac{3}{2}H^{2}(1-4\lambda),$ (40)
$c_{s}^{2}(\lambda)$ $=\frac{1}{3}-\frac{4\lambda}{3}.$ (41)
The Jeans wavenumber ($\omega=0$) is therefore:
$k_{J}^{2}=\frac{4\pi G_{\text{eff}}\rho}{c_{s}^{2}(\lambda)}a^{2}=\frac{\frac{3}{2}H^{2}(1-4\lambda)}{\frac{1}{3}(1-4\lambda)}a^{2}=\frac{9}{2}H^{2}a^{2}.$ (42)
Remarkably, the $\lambda$ dependence cancels exactly in the ratio, leaving the Jeans scale unchanged from GR. This is a nontrivial result of Rastall gravity during radiation domination—the modifications to gravity and the sound speed compensate each other at the linear level for the Jeans criterion.
5.1.2 Approach 2: Spherical collapse model
Adapting the spherical collapse formalism to Rastall gravity, the evolution equation for the scale factor $a_{p}$ of the perturbed region is derived from the modified Friedmann equations. For a top-hat overdensity $\delta$, the effective energy density inside the perturbed region is modified by the Rastall parameter:
$\rho_{\text{eff}}=\rho\left[1+\delta\left(1-\frac{3\lambda}{1-4\lambda}\right)\right].$ (43)
This leads to the modified collapse equation:
$\left(\frac{\dot{a}_{p}}{a_{p}}\right)^{2}=H^{2}\left[1+\delta\left(1-\frac{3\lambda}{1-4\lambda}\right)\right].$ (44)
The determination of the critical density contrast $\delta_{c}$ for primordial black hole formation requires a fully nonlinear treatment of gravitational collapse during radiation domination. In general relativity as mentioned before, $\delta_{c}^{\rm GR}\simeq 0.414$ has been obtained through relativistic numerical simulations that incorporate pressure gradients, profile dependence, and the compaction function formalism.
A rigorous determination of $\delta_{c}(\lambda)$ would require solving the fully nonlinear relativistic collapse equations in Rastall gravity, which lies beyond the scope of the present work. Instead, we adopt a phenomenological parameterization that captures the leading-order sensitivity of the collapse threshold to small deviations from general relativity. Restricting to the observationally allowed regime $|\lambda|\ll 1$, we expand the threshold as
$\delta_{c}(\lambda)=\delta_{c}^{\rm GR}\left(1+\gamma\lambda\right),$ (45)
where $\gamma$ is a dimensionless coefficient encoding the response of the nonlinear collapse dynamics to the Rastall modification.
The parameter $\gamma$ is expected to be $\mathcal{O}(1)$ in magnitude, but its precise value depends on the perturbation profile and on the detailed nonlinear evolution. In the absence of full numerical simulations, we treat $\gamma$ as a phenomenological parameter and explore representative values in the range $|\gamma|\lesssim 1$. Our conclusions regarding the exponential sensitivity of the PBH abundance to $\lambda$ are robust against moderate variations of $\gamma$.
A dedicated numerical study of relativistic collapse in Rastall gravity would be required to determine $\gamma$ from first principles. This positive $\gamma$ means that positive $\lambda$ makes collapse slightly harder (higher $\delta_{c}$), while negative $\lambda$ makes collapse easier.
Figure 2: Different profile dependences for critical density contrast.
The gray shaded region in Figure:1 represents the theoretical uncertainty in the critical density contrast arising from the dependence on the shape of the perturbation profile. Numerical simulations in general relativity [66,67] have established that $\delta_{c}$ varies by approximately $\pm 0.025$ depending on the specific profile shape (e.g., Gaussian, Mexican-hat, or top-hat profiles). In Rastall gravity, this uncertainty is expected to scale moderately with the modification parameter, as stronger deviations from general relativity amplify the profile dependence of gravitational collapse. In the Figure:2, we have shown the different profile dependences. Ofcourse just from the viewpoint of the critical density, increased critical density means more difficult to collapse. But as we will see due to the presence of $\lambda$ in the $\sigma$, the final abundance calculation gets different effect.
5.2 Variance of Density Fluctuations
The variance of density fluctuations smoothed on a comoving scale $R=1/k$ is:
$\sigma^{2}(R)=\int_{0}^{\infty}\frac{dk}{k}\mathcal{P}_{\delta}(k)W^{2}(kR),$ (46)
where $W(kR)$ is a window function (typically Gaussian or top-hat).
From the perturbation analysis, the density power spectrum at horizon crossing is related to the primordial curvature power spectrum by:
$\mathcal{P}_{\delta}(k)=\frac{16}{81}(1+2\beta\lambda)\mathcal{T}^{2}(k\tau)\mathcal{P}_{\mathcal{R}}(k),$ (47)
where the coefficient $\beta=\alpha_{H}$ arises from matching the super-horizon solution to the sub-horizon evolution at horizon crossing. and $\mathcal{T}(k\tau)$ is the transfer function.
For a scale-invariant primordial power spectrum $\mathcal{P}_{\mathcal{R}}(k)=A$, and smoothing on the horizon scale $R=1/(aH)$ at horizon crossing, the variance simplifies to:
$\sigma(M_{*})=\frac{4}{9}(1+\beta\lambda)\sqrt{A},$ (48)
where $M_{*}$ is the horizon mass at horizon crossing.
5.3 PBH Abundance
5.4 Primordial Black Hole Abundance
The abundance of primordial black holes forming at a given mass scale is computed using the Press–Schechter formalism under the assumption of Gaussian statistics. The mass fraction collapsing into PBHs at the time of formation is
$\beta(M)=\int_{\delta_{c}}^{\infty}\frac{d\delta}{\sqrt{2\pi}\,\sigma(M)}\exp\left(-\frac{\delta^{2}}{2\sigma^{2}(M)}\right)=\frac{1}{2}\mathrm{erfc}\left(\frac{\delta_{c}(\lambda)}{\sqrt{2}\,\sigma(M,\lambda)}\right),$ (49)
where $\sigma(M,\lambda)$ is the variance of density perturbations smoothed on the mass scale $M$, and $\delta_{c}(\lambda)$ is the critical threshold for collapse.
In the regime relevant for PBH formation, the variance is typically small, $\sigma\ll\delta_{c}$, and the abundance is dominated by the exponential tail of the Gaussian distribution. In this limit, Eq. (49) reduces to
$\beta(M)\approx\frac{\sigma(M,\lambda)}{\sqrt{2\pi}\,\delta_{c}(\lambda)}\exp\left(-\frac{\delta_{c}^{2}(\lambda)}{2\sigma^{2}(M,\lambda)}\right).$ (50)
We now incorporate the leading-order Rastall corrections. As discussed in the previous subsections, we parameterize the collapse threshold and the variance as
$\delta_{c}(\lambda)$ $=\delta_{c}^{\rm GR}\left(1+\gamma\lambda\right),$ (51)
$\sigma(M,\lambda)$ $=\sigma_{\rm GR}(M)\left(1+\beta\lambda\right),$ (52)
where $\gamma$ encodes the modification of the nonlinear collapse threshold, and $\beta$ arises from the modification of the density contrast at horizon crossing.
It is convenient to define the peak significance parameter
$\nu_{c}(\lambda)\equiv\frac{\delta_{c}(\lambda)}{\sigma(M,\lambda)}.$ (53)
Expanding to linear order in $\lambda$ (with $|\lambda|\ll 1$), we obtain
$\nu_{c}(\lambda)=\nu_{c}^{\rm GR}\left[1+(\gamma-\beta)\lambda\right],$ (54)
where $\nu_{c}^{\rm GR}=\delta_{c}^{\rm GR}/\sigma_{\rm GR}$.
Substituting Eq. (54) into Eq. (50) and keeping terms to leading order in $\lambda$ inside the exponential yields
$\frac{\beta(\lambda)}{\beta_{\rm GR}}\approx\exp\left[-(\nu_{c}^{\rm GR})^{2}(\gamma-\beta)\lambda\right],$ (55)
where $\beta_{\rm GR}$ denotes the abundance computed in general relativity.
Equation (55) demonstrates that even a small Rastall parameter $|\lambda|\lesssim 0.1$ can lead to exponentially large enhancements or suppressions of the PBH abundance, due to the factor $(\nu_{c}^{\rm GR})^{2}$ multiplying the correction. For typical rare-peak scenarios where $\nu_{c}^{\rm GR}\sim 5$1010, the exponential sensitivity implies that percent-level deviations in the collapse dynamics can modify the PBH abundance by several orders of magnitude.
Figure 3: Enhancement factor for different values of A.
6 Implications for PBH Dark Matter Scenarios
Observational constraints on primordial black hole (PBH) abundances from microlensing, cosmic microwave background (CMB) distortions, dynamical effects, and gravitational-wave observations leave several mass windows in which PBHs could constitute a significant fraction, or even the entirety, of dark matter. Representative examples include the asteroid-mass window ($10^{17}$$10^{22}$ g), the lunar-mass window ($10^{20}$$10^{24}$ g), and portions of the stellar and intermediate-mass range.
In general relativity, producing an appreciable PBH abundance in any of these windows requires a substantial enhancement of the primordial curvature power spectrum relative to its CMB value $\mathcal{P}_{\mathcal{R}}\simeq 2\times 10^{-9}$. Typically, amplitudes as large as $\mathcal{P}_{\mathcal{R}}\sim 10^{-2}$$10^{-1}$ are required at small scales, implying a degree of fine-tuning in inflationary model building.
In Rastall gravity, the background expansion during radiation domination remains identical to GR, but perturbation dynamics are modified at linear order in the Rastall parameter $\lambda$. As shown in the previous sections, the PBH abundance depends on the combination
$\nu_{c}(\lambda)=\nu_{c}^{\rm GR}\left[1+(\gamma-\beta)\lambda\right],$ (56)
where $\beta$ encodes the modification of the horizon-crossing amplitude and $\gamma$ parameterizes the shift in the nonlinear collapse threshold.
Since the abundance scales approximately as
$\beta(\lambda)\sim\exp\!\left[-\frac{\nu_{c}^{2}(\lambda)}{2}\right],$ (57)
even small deviations from GR can produce exponentially large changes in the PBH fraction. The sign and magnitude of the effect are governed by the difference $(\gamma-\beta)$:
  • If $(\gamma-\beta)\lambda<0$, the effective peak significance $\nu_{c}$ decreases and PBH formation is exponentially enhanced.
  • If $(\gamma-\beta)\lambda>0$, the effective peak significance increases and PBH formation is suppressed.
Thus, Rastall gravity does not universally enhance PBH formation. Rather, it modifies the abundance in a direction that depends on the relative size of the linear perturbation correction and the nonlinear collapse response.
For rare-peak scenarios with $\nu_{c}^{\rm GR}\sim 5$1010, percent-level deviations in $(\gamma-\beta)\lambda$ can shift the required primordial amplitude by $\mathcal{O}(10\%)$ while producing order-of-magnitude changes in the resulting PBH abundance. Consequently, even modest deviations from GR may significantly relax —or tighten—the fine-tuning requirements on the small-scale primordial power spectrum.
An important implication is that PBH constraints can be reinterpreted as constraints on the combination $(\gamma-\beta)\lambda$. Because the background cosmology during radiation domination is unchanged, PBH formation provides a particularly clean probe of perturbation-level modifications in Rastall gravity.
Finally, if future observations were to establish a PBH dark matter component with a well-measured mass function, it may be possible to distinguish Rastall gravity from GR through the inferred relation between the primordial power spectrum amplitude and the observed abundance. In this sense, PBHs offer a complementary avenue to large-scale structure and CMB tests for probing small departures from standard gravitational dynamics.
7 Conclusions and Outlook
In this work we have presented the first systematic analysis of primordial black hole (PBH) formation in Rastall gravity during radiation domination. A central result of our study is that, although the homogeneous background evolution remains identical to general relativity (GR) for a pure radiation fluid, the perturbation dynamics are modified at linear order in the Rastall parameter $\lambda$.
We derived the modified evolution equation for the density contrast during radiation domination and showed that Rastall gravity alters both the effective gravitational driving term and the effective sound speed. Interestingly, at linear order in $\lambda$, the Jeans scale remains unchanged due to a cancellation between these two effects. However, the super-horizon evolution and the amplitude of the density contrast at horizon crossing acquire $\lambda$-dependent corrections. As a consequence, the variance of density perturbations on the horizon scale is modified as
$\sigma(M,\lambda)=\sigma_{\rm GR}(M)\left(1+\beta\lambda\right),$ (58)
where $\beta$ is determined by the linear perturbation matching.
The nonlinear collapse threshold $\delta_{c}$ cannot be determined from linear theory alone. Since a full relativistic collapse analysis in Rastall gravity has not yet been performed, we adopted a phenomenological parameterization
$\delta_{c}(\lambda)=\delta_{c}^{\rm GR}\left(1+\gamma\lambda\right),$ (59)
where $\gamma$ encodes the response of nonlinear collapse dynamics to the modified gravitational coupling.
Combining these effects, we showed that the peak significance parameter becomes
$\nu_{c}(\lambda)=\nu_{c}^{\rm GR}\left[1+(\gamma-\beta)\lambda\right].$ (60)
Because the PBH abundance depends exponentially on $\nu_{c}^{2}$, even small deviations from GR can produce exponentially large enhancements or suppressions:
$\frac{\beta(\lambda)}{\beta_{\rm GR}}\sim\exp\!\left[-(\nu_{c}^{\rm GR})^{2}(\gamma-\beta)\lambda\right].$ (61)
Our analysis demonstrates that PBH formation provides a sensitive probe of modified gravity even in scenarios where the background cosmology is indistinguishable from GR. The exponential sensitivity of the abundance makes PBHs particularly powerful for constraining small deviations from standard gravitational dynamics.
Outlook
Several important extensions of this work remain.
First, a definitive determination of the collapse threshold $\delta_{c}(\lambda)$ requires solving the fully nonlinear relativistic collapse equations in Rastall gravity. Such a study would fix the parameter $\gamma$ from first principles and determine whether the modification enhances or suppresses collapse in realistic perturbation profiles.
Second, the linear perturbation analysis could be extended to include possible non-adiabatic modes arising from the non-conservation of the energy–momentum tensor. A gauge-invariant treatment of curvature perturbations in Rastall gravity would clarify the robustness of the horizon-crossing modification.
Third, it would be valuable to compute the induced stochastic gravitational wave background associated with enhanced scalar perturbations in this framework. Since the background expansion is unmodified during radiation domination, any observable deviation would arise purely from perturbation dynamics, providing a clean diagnostic of the theory.
Finally, embedding Rastall gravity within a consistent early-universe model (for example, during inflation or reheating) would allow a self-consistent calculation of the primordial curvature power spectrum and its possible $\lambda$-dependent features.
In summary, primordial black hole formation offers a novel and highly sensitive window into Rastall gravity. Even when the background cosmology mimics general relativity, perturbation-level effects can leave exponentially amplified imprints on the PBH abundance, making future observational constraints on PBHs a promising probe of modified gravitational dynamics.
Acknowledgments
The author would like to thank Prof. M. Sami and Dr. Debottam Nandi for insightful comments during the preparation of the manuscript. The work of MRG is supported by the Science and Engineering Research Board (SERB), DST, Government of India, under the Grant Agreement number CRG/2022/004120 (Core Research Grant). MRG would like to thank IUCAA, Pune, India, for their hospitality during the visit as an associate, where this work was initiated.

### v2

@@
 We investigate their production in Rastall gravity, a modification of general relativity that introduces a non-minimal coupling between matter and geometry through the non-conservation of the energy-momentum tensor.
-Analyzing cosmological perturbations during radiation domination, we demonstrate that the Rastall parameter fundamentally alters the collapse dynamics, modifying the growth of density fluctuations, the critical threshold for black hole formation, and the fluctuation amplitude at horizon crossing.
+Analyzing cosmological perturbations during radiation domination, we demonstrate that the Rastall parameter fundamentally alters the collapse dynamics, modifying the growth of density fluctuations, the critical threshold for black hole formation, and the fluctuation amplitude at horizon crossing if we consider the nearly conformal plasma.
 Current cosmological constraints from Big Bang Nucleosynthesis, the cosmic microwave background, and large-scale structure restrict the Rastall parameter to small values, yet within this allowed range PBH production can be altered by orders of magnitude compared to general relativity.
@@
 1 Introduction
-Primordial black holes (PBHs) have emerged as compelling dark matter candidates and unique probes of the early universe [1, 2, 3, 4, 5, 6].
+Primordial black holes (PBHs) have emerged as compelling dark matter candidates and unique probes of the early universe [1, 2, 3, 4, 5, 8, 6, 7, 8].
 Formed from the gravitational collapse of large-amplitude density perturbations during radiation domination, PBHs bridge the gap between early universe physics and observational astronomy, particularly through gravitational wave astronomy [9, 10, 11, 12, 13, 14, 15].
-The standard formation mechanism in general relativity (GR) requires specific enhancements of the primordial power spectrum, often necessitating fine-tuned inflationary dynamics [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
+The standard formation mechanism in general relativity (GR) requires specific enhancements of the primordial power spectrum, often necessitating fine-tuned inflationary dynamics [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
 Modified gravity theories offer alternative pathways to achieve the conditions necessary for PBH formation [30, 31, 32, 33, 34].
@@
 Consequently, the background evolution for a single radiation fluid is indistinguishable from GR.
-While the background dynamics for pure radiation are identical to GR, differences emerge at the perturbation level.
+While the background dynamics for pure radiation are identical to GR, differences emerge at the perturbation level if we consider the deviation from perfectly conformal plasma.
+The details are given in 2.1.
 This makes PBH formation during radiation domination an ideal probe of Rastall gravity, as any deviations from GR must arise from perturbations rather than background expansion.
+2.1 Activation of Rastall Corrections in a Nearly Conformal Plasma
+In the strict radiation-dominated limit with equation of state $p=\rho/3$, the energy–momentum tensor is traceless,
+$T\equiv T^{\mu}_{\ \mu}=-\rho+3p=0.$ (18)
+In this limit, the modified conservation law of Rastall gravity,
+$\nabla_{\mu}T^{\mu\nu}=\gamma\nabla^{\nu}R,$ (19)
+reduces identically to the standard conservation equation, and the theory becomes dynamically equivalent to General Relativity.
+Consequently, in an exactly conformal radiation fluid, Rastall corrections would be absent both at background and linear perturbation levels.
+However, the primordial plasma is not exactly conformal.
+Around the QCD epoch and at temperatures relevant for horizon re-entry of small-scale perturbations, quantum effects generate a small but non-vanishing trace of the energy–momentum tensor due to the QCD trace anomaly and deviations from ideal radiation behaviour.
+To capture this effect in a model-independent manner, we parametrize the deviation from conformality as
+$p=\left(\frac{1}{3}-\delta\right)\rho,\qquad\delta\ll 1,$ (20)
+which implies
+$T=-\rho+3p=-3\delta\rho.$ (21)
+Thus, while the background expansion remains effectively radiation-dominated, the trace is small but non-zero,
+$T\neq 0,$ (22)
+and the Rastall modification becomes active in a controlled manner.
+At linear order,
+$\delta T=-\delta\rho+3\delta p=-3\delta\,\delta\rho,$ (23)
+so that perturbatively the trace does not vanish.
+The modified conservation law therefore contributes at first order in $\delta$, leading to corrections proportional to the product $\gamma\delta$ in the scalar perturbation equation.
 3 Scalar Perturbations and Density Contrast Evolution
@@
 At horizon crossing, matching the super-horizon solution to the numerical solution gives the amplitude:
-$\delta_{H}(k)=\frac{4}{9}(1+\alpha_{H}\lambda)\mathcal{P}_{\mathcal{R}}^{1/2}(k),$ (39)
-where $\alpha_{H}\approx 0.3$ is a numerical coefficient derived from matching, and $\mathcal{P}_{\mathcal{R}}(k)$ is the primordial curvature power spectrum.
+$\delta_{H}(k)=\frac{4}{9}(1+\gamma\lambda)\mathcal{P}_{\mathcal{R}}^{1/2}(k),$ (45)
+where $\gamma$ is a numerical coefficient derived from matching, and $\mathcal{P}_{\mathcal{R}}(k)$ is the primordial curvature power spectrum.
 5 PBH Formation Criterion in Rastall Gravity
@@
 where $\gamma$ is a dimensionless coefficient encoding the response of the nonlinear collapse dynamics to the Rastall modification.
+Equation 51 should therefore be interpreted as the leading-order perturbative evolution equation in a nearly conformal plasma, where the effective deviation from GR is controlled by the small parameter
+$\gamma_{\rm eff}\equiv\gamma\delta.$ (52)
 The parameter $\gamma$ is expected to be $\mathcal{O}(1)$ in magnitude, but its precise value depends on the perturbation profile and on the detailed nonlinear evolution.
@@
 where $M_{*}$ is the horizon mass at horizon crossing.
-5.3 PBH Abundance
 5.3 Primordial Black Hole Abundance